OpenStudy (anonymous):

to solve the equation 5 sin(2x) = 3cosx you should write it as a. cosx (5sinx-3)=0 b. 5sinx+3cosx=0 c. cosx(10sin-3)=0

OpenStudy (anonymous):

\[ 5 \sin(2x) = 3\cos (x)\\ 10 \sin(x) \cos(x) = 3 \cos(x) \] Here we use the formula \[ \cos(2 a) = 2 \cos(a) \sin(a) \]

OpenStudy (anonymous):

Can you finish it now?

OpenStudy (anonymous):

\[ 5 \sin(2x) = 3\cos (x)\\ 10 \sin(x) \cos(x) = 3 \cos(x)\\ 10 \sin(x) \cos(x) - 3 \cos(x)=0\\ \cos(x)\left(10 \sin(x) - 3 \right)=0\\ \]